# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v2

@@ 4.2 Allowing Negative Eigenvalues
This modification causes that ${\bm{A}}^{-}_{\mathrm{diag}}({\bm{x}})$ has eigenvalues $2s({\bm{x}})_{i}-1\in[-1,1]$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})$ has all eigenvalues equal to one, except for one that is equal to $1-2\phi({\bm{x}})\in[-1,1]$. Thus, we have extended the range of eigenvalues from $[0,1]$ to $[-1,1]$.
@@
If we set the state-transition matrix in (1) to ${\bm{A}}(1)={\bm{R}}(\theta)$, an LRNN with one layer can count modulo $m$, since if we also set ${\bm{H}}_{0}=(1,0)^{\top}$ and $\mathrm{dec}({\bm{H}},x)=\arg\max_{i}{\bm{D}}_{i}^{\top}{\bm{H}}$, with ${\bm{D}}_{i}={\bm{R}}(i\theta){\bm{H}}_{0}$ for all $i\in\{0,\dots,m-1\}$, then for the input ${\bm{x}}=1^{t}$ we get
@@ 4.3 Expressivity of Products of Generalized Householder Matrices
We observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflections. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $\Omega\subseteq\Omega^{\prime}$ and either $k^{\prime}=k$ or $k^{\prime}\geq k$, $1\in\Omega$.
Our next result shows that products of GH matrices can represent any matrix with Euclidean norm less than or equal to 1. However, there are some restriction if we only allow each GH matrix in the product to have positive eigenvalues. In contrast, repeated products of upper (lower) triangular matrices with eigenvalues in $[-1,1]$ remain upper (lower) triangular, with eigenvalues in the same range.
@@
    Any eigenvalue $\lambda$ of ${\bm{N}}\in\mathcal{M}_{k}^{n}((-1,1])$ is either $1$ or satisfies $|\lambda|<1$ and if in addition ${\bm{N}}\in\mathcal{M}_{k}^{n}([0,1])$ and $k\leq 2$, then $\lambda\in[0,1]\subset\mathbb{R}$.
The proof in Section C.1 uses mainly linear algebra arguments such as the SVD decomposition and the fact that every $n\times n$ orthogonal matrix can be written as a product of $n$ reflections.
A consequence of Proposition 1 is that if for every layer of an LRNN, there exists $n,k\in\mathbb{N}$ such that the map ${\bm{A}}$ from inputs to state-transition matrices is such that ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([0,1])$, then if $k\leq 2$ the LRNN cannot learn parity, due to Theorem 1, and for any $k$ the state transition matrices are either the identity or not orthogonal, and hence unable to represent reflections or rotations. In contrast, if we allow ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([-1,1])$ and $k$ is large enough, the following theorem shows that an LRNN with one layer can implement any FSA whose transition monoid is a group, and that $n=k=2$ is enough for cyclic groups (modular addition).
@@
In the proof in Section C.2, we map each state-transition function to its matrix representation. This can always be done using permutation matrices, but for cyclic groups, we can also use rotation matrices. In the case of permutations, if every state-transitions permutes at most $k$ states then the corresponding permutation matrix will be in $\mathcal{M}_{k-1}^{n}(\{-1,1\})$, since it is either the identity or can be written as a product of at most $k-1$ permutations of two elements (swaps), each in $\mathcal{M}_{1}^{n}(\{-1\})$.
@@
The proof in Section C.4 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately. Note, that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards hybrids LRNNs using a mix of GH and diagonal matrices, whose exploration we leave for future work.
Discussion The results in Theorems 3 and 4 for LRNNs are in sharp contrast with the ones for Transformers (Liu et al., 2023; Merrill & Sabharwal, 2023) and diagonal LRNNs (Merrill et al., 2024), which require either the number of layers or the precision growing with the input sequence length, and can only implement an FSA if all groups in its transition monoid are solvable, i.e. excluding groups isomorphic to $S_{n}$ with $n\geq 5$. Moreover, compared to LRNNs without any restriction to the norm of the state-transition matrices, which need only one layer to recognize any regular language, our result requires both the number of layers and the width of the LRNN to be (in the worst case) exponential in the number of states of the FSA, although we conjecture that the number of layers might be reduced to at most linear using a more refined decomposition.
@@ 5.1 Chomsky Hierarchy
Parity Mod. Arithm. (w/o brackets) Mod. Arithm. w/ brackets)
@@
We conducted experiments with some of the formal language tasks proposed by Deletang et al. (2023) and similarly used to benchmark xLSTM (Beck et al., 2024). Our focus was on tasks where mLSTM (an LRNN) previously underperformed while sLSTM (a non-linear RNN) succeeded, specifically parity, modular arithmetic without brackets (both regular languages). and modular arithmetic with brackets (context-free language). As in Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to assess length generalization. Note that our theoretical results cover just regular languages, excluding modular arithmetic with brackets.
We compared a Transformer, mLSTM and sLSTM against two variants each of Mamba and DeltaNet - with and without eigenvalue range extension. Our findings, presented in Table 3, demonstrate that expanding the range of eigenvalues from $[0,1]$ to $[-1,1]$ enables all examined models to fully solve the parity task, confirming Theorem 1. For both modular arithmetic tasks, this expansion led to substantial performance improvements for Mamba and especially DeltaNet, since the latter has non-diagonal state-transition matrices that are more suited for these tasks (see Theorem 3). In Figure 4 in the Appendix, we visualize the length extrapolation performance of each model on all considered tasks. Note that we were unable to reproduce the sLSTM results reported by Beck et al. (2024) for the modular arithmetic tasks. Additional experiments and details on the tasks in Section E.1.
@@ 5.2 State-Tracking
We perform experiments on group word problems, relying on the code provided by Merrill et al., 2024. https://github.com/jopetty/word-problem. In particular, we focus on the $S_{5}$ group, which is the first unsolvable symmetric group where current LRNN and Transformers are known to perform poorly. We also report results for the addition modulo $60$, i.e. the cyclic group $\mathbb{Z}_{60}$, in Section E.2.2. We note that parity is $S_{2}$. In these experiments, the input to the model is a sequence of group elements, while the supervision is given by another sequence of group elements, each being the product of the previous ones in the input. Since solving $S_{5}$ would require LRNNs with state-transition matrices that are repeated products of 4 GH matrices (see Theorem 3), each with eigenvalues $[-1,1]$, we also consider three simplified setups: (i) allowing as inputs only permutations up to 2 elements (identity and swaps), (ii) allowing only permutations up to 3 elements, (iii) using 4 tokens for each permutation. Additional details are in Section E.2. We stress that, even when restricting the inputs, possible outputs remain the same, since swaps are generators of the group.
Results Figure 2 shows that, as predicted by Theorem 3, restricting the inputs to only swap permutations allows DeltaNet $[-1,1]$ with even one layer to fully learn the task (since its state-transition matrix can model a swap), while DeltaNet $[0,1]$ with 5 layers generalizes just slightly beyond the training length. In contrast, by including also permutations of $3$ elements, we notice a substantial decrease in the performance of all models. Interestingly, extending the range is still advantageous in this case and DeltaNet $[-1,1]$ with 5 layers reaches a good length generalization. Moreover, using 4 tokens per group element seems also beneficial compared to standard $S_{5}$, since DeltaNet $[-1,1]$ with 5 layers manages to extrapolate very well until around length $200$, which corresponds to $50$ group elements, while on standard $S_{5}$ all models have 00 sequence accuracy prior to sequence length $30$. We also report that Mamba, a diagonal LRNN, performs poorly on all setups, with and without increased eigenvalue range.
@@ 5.3 Language Modeling
Figure 3: Performance vs sequence length of DeltaNet variants (370M (top) and 1.3B (bottom) parameters) on four datasets. DeltaNet with eigenvalue range $[-1,1]$ improves perplexity in coding and math compared to the $[0,1]$ baseline. Dashed vertical line at training context length (2048).
@@
     Model Wiki. LMB. LMB. PIQA Hella. Wino. ARC-e ARC-c Avg. SWDE SQUAD FDA ppl $\downarrow$ ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ $\uparrow$ cont. $\uparrow$ cont. $\uparrow$ cont. $\uparrow$ 15B tokens SPJ 340M params        Transformer++ 28.39 42.69 31.0 63.3 34.0 50.4 44.5 24.2 41.2 42.2 22.1 21.4        Mamba $[0,1]$ 28.39 39.66 30.6 65.0 35.4 50.1 46.3 23.6 41.8 12.4 23.0 2.1        GLA $[0,1]$ 29.47 45.53 31.3 65.1 33.8 51.6 44.4 24.6 41.8 24.0 24.7 7.3        DeltaNet $[0,1]$ 28.24 37.37 32.1 64.8 34.3 52.2 45.8 23.5 42.1 26.4 28.9 12.8 100B tokens FW 340M params        DeltaNet $[0,1]$ 24.68 31.49 33.7 70.3 45.1 51.3 50.0 26.1 46.1 35.2 28.7 11.8        DeltaNet $[-1,1]$ 24.54 31.15 34.0 69.9 44.6 51.9 50.0 24.4 45.8 37.2 33.1 6.6 370M params        Mamba $[0,1]$ 24.84 24.69 35.6 70.6 48.4 51.2 53.4 24.8 47.3 21.6 27.7 2.8        Mamba $[-1,1]$ 25.02 24.71 36.2 70.5 47.8 53.3 54.7 26.7 48.2 20.9 24.8 2.5 100B tokens SPJ 1.3B params      Transformer++ 16.85 13.44 48.9 70.8 49.6 53.6 56.0 26.5 50.9 66.6 31.5 27.4      Mamba $[0,1]$ 17.06 13.89 46.2 72.2 40.1 54.1 59.0 28.2 50.0 41.4 35.2 6.2      GLA $[0,1]$ 17.22 14.47 46.9 71.8 49.8 53.9 57.2 26.6 51.0 50.6 42.6 19.9      DeltaNet $[0,1]$ 16.87 12.21 48.9 71.2 50.2 53.6 57.2 28.3 51.6 49.5 37.4 17.2 100B T FW 1.3B params      DeltaNet $[0,1]$ 18.54 14.32 43.5 73.7 56.2 56.9 58.2 29.9 53.1 49.1 35.1 8.6      DeltaNet $[-1,1]$ 18.57 12.73 43.7 73.3 55.8 56.8 56.9 27.9 52.4 48.8 33.9 12.3
@@ 6 Conclusion
Limitations and Future work  Our modification is not directly compatible with a numerical technique used by some diagonal LRNNs such as Mamba2, GLA and mLSTM. In particular, these models rely on positive state-transition matrices to compute cumulative products in log space, which improves numerical accuracy and potentially training stability (see Section E.4 for details). Further research is needed to assess the impact of training large-scale language models with state-tracking capabilities. To this end, we aim to understand the potential downsides of increased expressivity. For example, we hypothesize a fundamental trade-off between state-tracking and associative recall, which is also of theoretical interest and could guide hybrid model design. Moreover, the theoretical expressivity of DeltaNet $[-1,1]$ with multiple layers is still unclear. We showed that it can solve addition modulo $m$ (in Appendix D) which is equivalent to the $\mathbb{Z}_{3}$ group word problem, but we do not know if it can also solve word problems for the symmetric groups $\mathbb{S}_{n}$ with $n\geq 3$.

### v3

@@ 4.2 Allowing Negative Eigenvalues
Hence, ${\bm{A}}^{-}_{\mathrm{diag}}({\bm{x}})$ has eigenvalues $2s({\bm{x}})_{i}-1\in[-1,1]$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})$ has one eigenvalue equal to $1-2\phi({\bm{x}})\in[-1,1]$. Thus, we have extended the eigenvalues range from $[0,1]$ to $[-1,1]$. The norm of the matrix is still less than or equal to one, keeping the recurrence stable at long sequence lengths.
@@
If we set the state-transition matrix in (1) to ${\bm{A}}(1)={\bm{R}}(\theta)$, an LRNN with one layer can count modulo $m$, since if we also set ${\bm{H}}_{0}=(1,0)^{\top}$ and $\mathrm{dec}({\bm{H}},x)=\arg\max_{i}{\bm{D}}_{i}^{\top}{\bm{H}}$, with ${\bm{D}}_{i}={\bm{R}}(i\theta){\bm{H}}_{0}$ for all $i\in\{0,\dots,m-1\}$, then for the input ${\bm{x}}=1^{t}$ and since ${\bm{R}}$ has period $2\pi$, we get
@@ 4.3 Expressivity of Products of Generalized Householder Matrices
Intuitively, higher $k$ means higher expressivity but also higher cost for matrix-vector products. Furthermore, as long as $\Omega\subseteq[-1,1]$, the norm of the matrices is bounded by one. We observe that if ${\bm{M}}\in\mathcal{M}_{1}^{n}(\{-1\})$, then ${\bm{M}}$ is a reflection (or Householder) matrix, and that for any ${\bm{x}}\in\mathbb{R}^{l}$, ${\bm{A}}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([0,1])$ and ${\bm{A}}^{-}_{\mathrm{GH}}({\bm{x}})\in\mathcal{M}_{1}^{n}([-1,1])$ so that with our change we also include reflections. Moreover, $\mathcal{M}_{k}^{n}(\Omega)\subseteq\mathcal{M}_{k^{\prime}}^{n}(\Omega^{\prime})$ if $\Omega\subseteq\Omega^{\prime}$ and either $k^{\prime}=k$ or $k^{\prime}\geq k$, $1\in\Omega$.
Our next result shows that products of GH matrices can represent any matrix with Euclidean norm less than or equal to 1, but only when $[-1,1]\subseteq\Omega$. In contrast, repeated products of (e.g. upper) triangular matrices with eigenvalues in $[-1,1]$ remain triangular, with eigenvalues in the same range.
@@
    Any eigenvalue $\lambda$ of any matrix ${\bm{N}}\in\mathcal{M}_{k}^{n}((-1,1])$ is either $1$ or satisfies $|\lambda|<1$ and if in addition ${\bm{N}}\in\mathcal{M}_{k}^{n}([0,1])$ and $k\leq 2$, then $\lambda\in[0,1]\subset\mathbb{R}$.
The proof in Section C.1 uses mainly linear algebra arguments such as the SVD decomposition and the fact that every $n\times n$ orthogonal matrix can be written as a product of $n$ reflections, due to the Cartan–Dieudonné Theorem (Gallier & Gallier, 2011).
A consequence of Proposition 1.3 is that LRNNs with layers of the form (1), where ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([0,1])$, have state transition matrices that are either the identity or not orthogonal, and hence cannot be reflections or rotations. Also, if $k\leq 2$ the eigenvalues are positive and hence the LRNN cannot learn parity due to Theorem 1. In contrast, if we allow ${\bm{A}}:\mathbb{R}^{l}\to\mathcal{M}_{k}^{n}([-1,1])$ and $k$ is large enough, the following theorem shows that an LRNN with one layer can implement any FSA whose transition monoid is a group, and that $n=k=2$ is enough for cyclic groups (modular addition).
@@
In the proof in Section C.2, we map each state-transition function to a matrix representation. This can always be done using permutation matrices, but for cyclic groups, we can also use rotation matrices. In the case of permutations, if every state-transition permutes at most $k$ states then the corresponding permutation matrix will be in $\mathcal{M}_{k-1}^{n}(\{-1,1\})$, since it is either the identity or can be written as a product of at most $k-1$ permutations of two elements (swaps), each in $\mathcal{M}_{1}^{n}(\{-1\})$.
@@
The proof in Section C.4 exploits the landmark Theorem by Krohn & Rhodes (1965), which states that every FSA can be decomposed as a cascade of simpler FSAs whose state-transition functions are either one-to-one or constant. Each layer of the LRNN will implement one FSA (with $n$ states) of the cascade using $n\times n$ permutation matrices, which are in $\mathcal{M}_{n-1}^{n}(\{-1,1\})$, for the one-to-one transitions, while for constant (state-independent) transitions it will set the corresponding state-transition matrix to $0\in\mathcal{M}_{n}^{n}(\{0\})$ and the function ${\bm{B}}$ appropriately. Note, that we can obtain the zero matrix only inefficiently as a product of $n$ GH matrices, while it could also be obtained with a single diagonal matrix. This points towards LRNNs using a mix of GH and diagonal matrices, as recently explored by Gated DeltaNet (Yang et al., 2025) and RWKV-7.
Discussion The results in Theorems 3 and 4 for LRNNs are in sharp contrast with the ones for Transformers (Liu et al., 2023; Merrill & Sabharwal, 2023) and diagonal LRNNs (Merrill et al., 2024), which require either the number of layers or the precision growing with the input sequence length, and can only implement an FSA if all groups in its transition monoid are solvable, i.e. excluding groups isomorphic to $S_{n}$ with $n\geq 5$. However, compared to LRNNs without any restriction to the norm of the state-transition matrices, which need only one layer to recognize any regular language, our result requires both the number of layers and the width of the LRNN to be (in the worst case) exponential in the number of states of the FSA, although we conjecture that the number of layers might be reduced to at most linear using a more refined decomposition.
@@ 5.1 Chomsky Hierarchy
Parity Mod. Arithm. (w/o brackets) Mod. Arithm. (w/ brackets)
@@
We conducted experiments with some of the formal language tasks proposed by Deletang et al. (2023) and similarly used to benchmark xLSTM (Beck et al., 2024). Our focus was on tasks where mLSTM (an LRNN) previously underperformed while sLSTM (a non-linear RNN) succeeded, specifically parity, modular arithmetic without brackets (both regular languages). and modular arithmetic with brackets (context-free language). As in Beck et al. (2024), we trained each model with sequence lengths ranging from 3 to 40 and evaluated on lengths from 40 to 256, to assess length generalization. Note that our theoretical results cover just regular languages, excluding modular arithmetic with brackets. We compared a Transformer, mLSTM and sLSTM against two variants each of Mamba and DeltaNet - with and without eigenvalue range extension.
Results Our findings, presented in Table 3, demonstrate that expanding the range of eigenvalues from $[0,1]$ to $[-1,1]$ enables all examined models to fully solve the parity task, confirming Theorem 1. For both modular arithmetic tasks, this expansion led to substantial performance improvements for Mamba and especially DeltaNet, since the latter has non-diagonal state-transition matrices that are more suited for these tasks (see Theorem 3). In Figure 4 in the Appendix, we visualize the length extrapolation performance of each model on all considered tasks. Note that we were unable to reproduce the sLSTM results reported by Beck et al. (2024) for the modular arithmetic tasks. Additional experiments and details on the tasks in Section E.1.
@@ 5.2 State-Tracking
We perform experiments on group word problems, relying on the code provided by Merrill et al., 2024. We focus on the $S_{5}$ group—the first unsolvable symmetric group where current LRNNs and Transformers are known to underperform. We also report results for addition modulo $60$ (i.e., the cyclic group $\mathbb{Z}_{60}$) in Section E.2.2, and note that parity corresponds to $S_{2}$. In these experiments, the model receives a sequence of group elements as input, and the supervision is another sequence of group elements, each representing the product of the preceding input elements. Since solving $S_{5}$ might need LRNNs with state-transition matrices formed by repeated products of four GH matrices (see Theorem 3), each with eigenvalues in $[-1,1]$, we also consider three simplified setups: (i) allowing only permutations of up to 2 elements (identity and swaps), (ii) allowing only permutations of up to 3 elements, and (iii) using 4 tokens for each permutation. Additional details are in Section E.2. We stress that, even when restricting the inputs to only identity and swaps, the group elements for the supervision still cover the entire group, because swaps are generators of the group.
Results Figure 2 shows that, as predicted by Theorem 3, restricting the inputs to only swap permutations allows DeltaNet $[-1,1]$ with even one layer to fully learn the task (since its state-transition matrices can model swaps), while DeltaNet $[0,1]$ with 5 layers generalizes just slightly beyond the training length. In contrast, by including also permutations of $3$ elements, we notice a substantial decrease in the performance of all models. Interestingly, extending the range is still advantageous in this case and DeltaNet $[-1,1]$ with 5 layers reaches a good length generalization. Moreover, using 4 tokens per group element seems also beneficial compared to standard $S_{5}$, since DeltaNet $[-1,1]$ with 5 layers manages to extrapolate very well until around length $200$, which corresponds to $50$ group elements, while on standard $S_{5}$ all models have 00 sequence accuracy prior to sequence length $30$. We also report that Mamba, a diagonal LRNN, performs poorly on all setups, with and without increased eigenvalue range.
@@ 5.3 Language Modeling
Figure 3: Performance vs sequence length of DeltaNet variants (340M (top) and 1.3B (bottom) parameters) on four datasets. DeltaNet with eigenvalue range $[-1,1]$ improves perplexity in coding and math compared to the $[0,1]$ baseline. Dashed vertical line at training context length (2048).
@@
     Model Wiki. LMB. LMB. PIQA Hella. Wino. ARC-e ARC-c Avg. SWDE SQUAD FDA ppl $\downarrow$ ppl $\downarrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc_n $\uparrow$ $\uparrow$ cont. $\uparrow$ cont. $\uparrow$ cont. $\uparrow$ SlimPajama 15B 340M params        Transformer++ 28.39 42.69 31.0 63.3 34.0 50.4 44.5 24.2 41.2 42.2 22.1 21.4        Mamba $[0,1]$ 28.39 39.66 30.6 65.0 35.4 50.1 46.3 23.6 41.8 12.4 23.0 2.1        GLA $[0,1]$ 29.47 45.53 31.3 65.1 33.8 51.6 44.4 24.6 41.8 24.0 24.7 7.3        DeltaNet $[0,1]$ 28.24 37.37 32.1 64.8 34.3 52.2 45.8 23.5 42.1 26.4 28.9 12.8 FineWeb 100B 340M params        DeltaNet $[0,1]$ 24.68 31.49 33.7 70.3 45.1 51.3 50.0 26.1 46.1 35.2 28.7 11.8        DeltaNet $[-1,1]$ 24.54 31.15 34.0 69.9 44.6 51.9 50.0 24.4 45.8 37.2 33.1 6.6 370M params        Mamba $[0,1]$ 24.84 24.69 35.6 70.6 48.4 51.2 53.4 24.8 47.3 21.6 27.7 2.8        Mamba $[-1,1]$ 25.02 24.71 36.2 70.5 47.8 53.3 54.7 26.7 48.2 20.9 24.8 2.5 SlimPajama 100B 1.3B params      Transformer++ 16.85 13.44 48.9 70.8 49.6 53.6 56.0 26.5 50.9 66.6 31.5 27.4      Mamba $[0,1]$ 17.06 13.89 46.2 72.2 40.1 54.1 59.0 28.2 50.0 41.4 35.2 6.2      GLA $[0,1]$ 17.22 14.47 46.9 71.8 49.8 53.9 57.2 26.6 51.0 50.6 42.6 19.9      DeltaNet $[0,1]$ 16.87 12.21 48.9 71.2 50.2 53.6 57.2 28.3 51.6 49.5 37.4 17.2 FW 100B 1.3B params      DeltaNet $[0,1]$ 18.54 14.32 43.5 73.7 56.2 56.9 58.2 29.9 53.1 49.1 35.1 8.6      DeltaNet $[-1,1]$ 18.57 12.73 43.7 73.3 55.8 56.8 56.9 27.9 52.4 48.8 33.9 12.3
@@ 6 Conclusion
Limitations and Future work  Our modification is not directly compatible with a numerical technique used by some diagonal LRNNs such as Mamba2, GLA and mLSTM. In particular, these models rely on positive state-transition matrices to compute cumulative products in log space, which improves numerical accuracy and potentially training stability (see Section E.4 for details). Further research is needed to assess the impact of training large-scale language models with state-tracking capabilities. To this end, we aim to understand the potential downsides of increased expressivity. For example, we hypothesize a fundamental trade-off between state-tracking and associative recall, which is also of theoretical interest and could guide hybrid model design. Moreover, the theoretical expressivity of DeltaNet $[-1,1]$ with multiple layers is still unclear. We showed that it can solve addition modulo $m$ (in Appendix D) which is equivalent to the $\mathbb{Z}_{3}$ group word problem, but we do not know if it can also solve other word problems, such as the ones for the symmetric groups $S_{n}$ with $n\geq 3$.